# Quantum Support Vector Regression for Robust Anomaly Detection

### v2

@@ III-D Generation of Adversarial Attacks
IV Results and Discussion
In this study, we first benchmarked our QSVR on the 27-qubit IBM Ehningen device (labeled qc-QSVR) and compared its performance against the simulated quantum baseline models QSVR (simulated version of our model, see [Tscharke]) and QAE (based on [Kottmann2021]), as well as CSVR and CAE as classical baselines. Second, six different noise channels of varying strength were introduced to evaluate the influence of noise on the QSVR algorithm. Third, the adversarial robustness of the model was examined, and the influence of noise on the adversarial robustness was evaluated by exposing the (noisy) models to adversarial attacks. The simulations show no error bars because the SVR is a deterministic model and pennylane’s default.mixed [^2] https://docs.pennylane.ai/en/stable/code/api/pennylane.devices.default_mixed.DefaultMixed.html device used to calculate the kernels computes exact outputs.
IV-A Hardware Results
Figure 2 compares model performance on eleven datasets using area under the ROC curve (AUC), a commonly used metric in AD that measures the trade-off between the true and false positive rates independent of a detection threshold. An AUC of 1.0 indicates perfect classification of the dataset, and a random classifier achieves an AUC of 0.5 on balanced binary datasets. The datasets include Credit Card Fraud (CC), Census, Forest Cover Type (CoverT), Domain Name System over HTTPS (DoH), EMNIST, Fashion MNIST (FMNIST), Network Intrusion (KDD), MNIST, Mammography (Mammo), URL, and our constructed dataset Toy. The models included in the study are qc-QSVR (ours), QSVR, QAE, CSVR, and CAE. The average performance of each model across all datasets is represented by a dotted line, while the bars indicate the models’ performance on individual datasets.
On average, our qc-QSVR exhibits an AUC decrease of 0.04 compared to the simulated QSVR. In 8 out of 11 datasets, the simulated model outperforms the qc-QSVR, with the performance gap explained by hardware noise. On the DoH datasets, both models perform identically, while on the CC and KDD datasets, the qc-QSVR surprisingly outperforms its simulated counterpart. On these two datasets, hardware-induced noise appears to enhance model performance, an effect we attribute to improved generalization. Specifically, the perturbations introduced by the noisy gates mimic the corruptions applied in denoising autoencoders, a technique shown to yield superior generalization compared to standard autoencoders [denoising_autoencoder]. Our results are consistent with those of other authors who observe that noise can improve the performance of QML models under certain circumstances [Escudero_2023, Domingo2023].
@@ IV-A Hardware Results
IV-B Influence of Noise on the Model Performance
The hardware results show that noise in NISQ devices affects the performance of QSVR. Therefore, in this section, we examine the influence of six noise channels on the QSVR based on the performance on eleven datasets. Figure 3 shows the AUC of noisy simulations with the noise channels described in Section II-B as well as the influence of adversarial attacks of strength $\varepsilon=0.1$ on the noisy simulations. The model’s robustness against noise is now analyzed, and the adversarial robustness is investigated in Section IV-C. The QSVR is largely robust against depolarizing, phase damping, phaseflip , and bitflip noise, as the AUC for these noise types remains rather stable with increasing noise probability. A theoretical analysis of the robustness of quantum classifiers done by LaRose and Coyle [noise_robustness_LaRose] proves that single-qubit classifiers are robust against precisely these noise channels. Taking into account the findings of Schuld [supervised_qml_are_kernel_schuld2021] that many short-term and fault-tolerant quantum models can be replaced by a general support vector machine whose kernel computes distances between data-encoding quantum states, we expect these results to transfer to our QSVR.
Amplitude damping has a large effect on the model’s behavior as the AUC decreases for all datasets except CoverT and URL at $p=0.1$ and $p=0.2$. At higher $p$, the AUC partially recovers, approaching 0.5. This shows that the QSVR is very sensitive to amplitude damping , and at high noise levels the model becomes a random classifier. These results are also supported by LaRose and Coyle’s [noise_robustness_LaRose] analysis, who found that quantum classifiers are generally not robust against amplitude damping noise.
The curves of the models subject to miscalibration noise show a periodicity of approximately $\pi$, with dips at about $k\pi$ for $k=0,1,2$, and plateaus in between. Small levels of miscalibration degrade the performance of the model, but when the noise level is above about $0.25\pi$ , the AUC reaches a plateau at a level similar to the one for zero noise until the curve dips again around $p=\pi$. We conclude that small degrees of miscalibration reduce model performance, and that this type of noise should be avoided in hardware.
@@ IV-C Adversarial Robustness
AD is often used in security-critical areas such as credit card fraud detection or network intrusion detection. Therefore, ML models used for AD must be robust to adversarial attacks. For low-dimensional, tabular data sets, it is possible that a sample can be completely and effectively transformed into a sample of a different class at high attack strengths. In these cases, the effect on the AUC may be exaggerated and should be interpreted only as an upper bound on the performance drop.
IV-C1 Noise-Free Adversarial Attacks
@@
First, we consider the noise-free model and plot the obtained results of the PGD attacks up to a strength of $\varepsilon=0.5$ in Figure 4. The noise-free QSVR is highly vulnerable to adversarial attacks, as evidenced by the decrease in AUC for small attack strengths of $\varepsilon=0.01$ for all datasets except Toy. The largest decrease is in the AUC of the DoH dataset, which drops by an order of magnitude from 0.67 to 0.06 for the $\varepsilon=0.01$ attack. As the attack strength increases, the AUC continues to decrease for all datasets until it approaches 0.0 for $p=0.3$. We conclude that techniques to increase the adversarial robustness of the model should be investigated.
@@
The unchanged AUC of Toy for $\varepsilon=0.01$ can be explained by the creation process of the dataset. The dataset was created to be linearly separable with a separation distance of 0.4 between the classes, so the data points must be shifted a large distance in feature space to be misclassified. However, since the AUC drops to 0.0 for $\varepsilon=0.3$, which is smaller than the separation distance, we might conclude that the QSVR is susceptible to overfitting.
IV-C2 Noisy Adversarial Attacks
@@
Second, the performance of the noisy QSVR when subjected to adversarial perturbations of strength $\varepsilon=0.1$ is shown in Figure 3. Omitting miscalibration, we find that for noise levels below $p=0.1$, the AUC is low for most noise types and datasets, following the trend of high adversarial vulnerability observed in Figure 4. At higher noise levels, however, the AUC typically recovers to some extent, reaching a plateau at about an AUC of 0.5. This indicates that the model transitions to a random classifier, showing that the adversarial attacks are so powerful that quantum noise cannot improve performance beyond that of a random classifier. Other researchers report similar results, noting that random noise and adversarial noise are fundamentally different, and that models resilient to random noise are often still vulnerable to adversarial noise [fawzi2016robustnessclassifiersadversarialrandom].
Miscalibration noise affects both models under attack and models not under attack, similarly, resulting in spikes around $p=k\pi$ for $k=0,1,2$ in most datasets. Interestingly, for adversarially attacked models, as opposed to models that are not under attack, these spikes can be lower than the adjacent plateaus, depending on the dataset. Between these spikes, the AUC generally remains stable, forming plateau regions.
@@
The DoH dataset is an outlier, with an AUC of 0.0 across almost all noise types and strengths, which is explained by its extreme vulnerability to adversarial attacks seen in Figure 4. This vulnerability can be explained by Figure 5 in Appendix B in the appendix, showing the p-value from the Kolmogorov–Smirnov test and the maximum feature variance within the test set. The DoH dataset has a relatively high p-value combined with a very low variance. The high p-value indicates a high probability that the normal and anomalous samples originate from the same distribution, while the low variance suggests a high degree of similarity between all samples, especially between the normal and anomalous data. As a result, the DoH dataset is difficult to classify, and even tiny adversarial attacks of strength $\varepsilon=0.01$ lead to manipulations a magnitude greater than the variance within the dataset.
@@
Notably, the AUC for the DoH dataset with miscalibration noise is $0.0$ over nearly all noise levels and peaks only around $p=\pi$, where it approaches $1.0$. An analysis of the adversarial test kernels for the DoH dataset subject to miscalibration noise is shown in Table III and highlights major differences between a high-performing run ($p=2.9\approx 0.9\pi$, $\text{AUC}=0.98$) and a low-performing run ($p=1.7\approx 0.5\pi$, $\text{AUC}=0.00$) for both classes. The mean kernel values for the high-AUC run are four orders of magnitude larger than those for the low-AUC run. In addition, the disparity between kernel values of classes 0 and 1 is greater in the high-performing scenario, allowing for easier distinction by the SVR and thus improved model AUC. Considering that the kernel entries represent the probability of measuring the all-zero state, we observe that miscalibration noise with a strength close to $\pi$ shifts the DoH-embedding states closer to the all-zero state, thus increasing the kernel values.
@@
Since the overrotation introduced by miscalibration noise is independent of the data, this type of noise can be thought of as additional fixed parameter rotation gates in the circuit. Because the parameters of these gates have a significant impact on model performance, we highlight the importance of using kernels tailored to the dataset, such as trainable kernels.
@@
TABLE III: Analysis of the adversarial test kernel for the DoH dataset subject to miscalibration noise
@@
IV-C3 Adversarial Training
Adversarial training is a straightforward approach to increasing the adversarial robustness of supervised learning algorithms. Table IV reveals that adversarial training increases the AUC for the adversarial test set on seven out of eleven datasets, and the average AUC over all datasets rises from 0.28 to 0.31. However, the increase in AUC is small, and except for FMNIST and Toy, the AUC remains below 0.5. For the test set without adversarial samples, the AUC decreases through adversarial training on six out of eleven datasets, and the average declines from 0.75 to 0.71. Table V shows the ratio of correctly classified normal samples to the total number of normal samples, as well as the same ratio for the anomalies. For normal data, the ratio is $\frac{\text{tn}}{\text{tn}+\text{fp}}$ , and for the anomalies it is $\frac{\text{tp}}{\text{tp}+\text{fn}}$. We observe that retraining increases the classification ratio for the normal data of the Toy dataset from 0.78 to 0.94, while the ratio for the anomalies remains unchanged at 1.00. For KDD, we report similar results, but the increase in the classification ratio of the normal data through retraining is smaller. This shows that adversarial retraining can lead to more normal samples being classified correctly without influencing the classification of the anomalies, since the latter are not contained in the training set. However, this was not observed for other datasets and was most pronounced for the synthetic dataset, suggesting this behavior requires a large separation distance between the two classes.
@@ IV-C3 Adversarial Training
TABLE IV: AUCs of the models for the test set and adversarial test set with and without adversarial training.
@@
TABLE V: Ratios of correctly classified normal and anomalous samples.
@@ V Conclusion and Outlook
Second, the influence of six noise channels on the performance of the QSVR was evaluated, revealing that the QSVR is largely robust against deporarizing, phasedamping, phase flip and bit flip noise. Amplitude damping, on the other hand, results in the most significant degradation of the model and miscalibration noise also has the potential to impact performance.
Finally, the adversarial robustness of the (noisy) model was assessed, and it was observed that the QSVR is highly vulnerable to adversarial attacks. Even weak PGD attacks with a strength of $\varepsilon=0.01$ can reduce the AUC by up to an order of magnitude. Introducing quantum noise does not yield any beneficial effect, neither on the unattacked model’s performance nor on its adversarial robustness. Moreover, adversarial training does not reliably improve the adversarial robustness of the model. Consequently, we conclude that the QSVR demonstrates potential for semisupervised AD in the NISQ era, however, special attention should be paid to the vulnerability to adversarial attacks and amplitude damping and miscalibration noise.
We emphasize the importance of employing dataset-specific kernels and recommend exploring trainable kernels to further enhance the performance of QML models. Future research directions could also include expanding the model to more qubits, and finally, future work may also focus on techniques to enhance the adversarial robustness of the QSVR and defend against such attacks.
Acknowledgment
@@
[…]
Appendix A Overview of the Datasets
TABLE VI: Overview of the datasets used for the experiments.
@@ Appendix A Overview of the Datasets
CC [cc] Normal Anomalous
Census [census_income] $\leq$50k $>$50k

### v3

@@ 3.4 Generation of Adversarial Attacks
4 RESULTS AND DISCUSSION
In this study, we first benchmarked our QSVR on the 27-qubit IBM Ehningen device (labeled qc-QSVR) and compared its performance against the simulated quantum baseline models QSVR (simulated version of our model, see [Tscharke]) and QAE (based on [Kottmann2021]), as well as CSVR and CAE as classical baselines. Second, six different noise channels of varying strength were introduced to evaluate the influence of noise on the QSVR algorithm. Third, the adversarial robustness of the model was examined, and the influence of noise on the adversarial robustness was evaluated by exposing the (noisy) model to adversarial attacks. The simulations show no error bars because the SVR is a deterministic model and pennylane’s default.mixed [^2] https://docs.pennylane.ai/en/stable/code/api/pennylane.devices.default˙mixed.DefaultMixed.html device used to calculate the kernels computes exact outputs.
4.1 Hardware Results
Figure 2 compares model performance on eleven datasets using area under the ROC curve (AUC). The datasets include Credit Card Fraud (CC), Census, Forest Cover Type (CoverT), Domain Name System over HTTPS (DoH), EMNIST, Fashion MNIST (FMNIST), Network Intrusion (KDD), MNIST, Mammography (Mammo), URL, and our constructed dataset Toy. The models included in the study are qc-QSVR (ours), QSVR, QAE, CSVR, and CAE. The average performance of each model across all datasets is represented by a dotted line, while the bars indicate the models’ performance on individual datasets.
On average, our qc-QSVR exhibits an AUC decrease of 0.04 compared to the simulated QSVR. In 8 out of 11 datasets, the hardware model is outperformed by the simulated version, with the performance gap attributed to hardware noise. On the DoH datasets, both models perform identically, while on the CC and KDD datasets, the qc-QSVR surprisingly outperforms its simulated counterpart. On these two datasets, hardware-induced noise appears to enhance model performance, an effect we ascribe to improved generalization. Specifically, the perturbations introduced by the noisy gates might mimic the corruptions applied in denoising autoencoders, a technique shown to yield superior generalization compared to standard autoencoders [denoising_autoencoder]. These results are consistent with those of other authors showing that noise can, under certain conditions, improve the performance of QML models [Escudero_2023, Domingo2023]. This phenomenon of superior model performance caused by hardware noise will be the subject of future research.
@@ 4.1 Hardware Results
4.2 Influence of Noise on the Model Performance
The hardware results show that noise in NISQ devices usually affects the results obtained by the QSVR. Therefore, we examine the influence of six noise channels on the performance of the QSVR on eleven datasets in this section. Figure 3 presents the AUC of simulations across different noise levels for the channels described in Section 2.2, along with the impact of adversarial attacks of strength $\varepsilon=0.1$ on these noisy simulations. The influence of noise on the adversarial robustness is investigated later, in Section 4.3. The QSVR is largely robust against depolarizing, phase damping, phase flip, and bit flip noise, as the AUC for these noise types remains rather stable with increasing noise probability. A theoretical analysis of the robustness of quantum classifiers done by [noise_robustness_LaRose] proves that single-qubit classifiers are robust against precisely these noise channels. Taking into account the findings of [supervised_qml_are_kernel_schuld2021] that many short-term and fault-tolerant quantum models can be replaced by a general support vector machine whose kernel computes distances between data-encoding quantum states, we expect these results to transfer to our QSVR.
Amplitude damping has a large effect on the model’s behavior as the AUC decreases for all datasets at $p=0.1$ and $p=0.2$ except for CoverT and URL. At higher $p$, the AUC partially recovers and approaches roughly 0.5. These findings indicate that the QSVR is highly sensitive to amplitude damping, with performance under high noise degrading to that of a random classifier. Our observation is consistent with the analysis of [noise_robustness_LaRose], who showed that quantum classifiers lack robustness against this type of noise.
The curves of the models subject to miscalibration noise show a periodicity of approximately $\pi$, with dips at about $k\pi$ for $k=0,1,2$, and plateaus in between. Small levels of miscalibration degrade the performance of the model, but when the noise level is above about $0.25\pi$ , the AUC reaches a plateau at a level similar to the one for zero noise until the curve dips again around $p=\pi$. The miscalibration introduces additional rotations in the circuit, effectively changing the states after feature encoding. This alters the fidelities and thereby impacts the reconstruction error and overall performance. We conclude that small degrees of miscalibration reduce model performance, and that this type of noise should be avoided in hardware. However, we leave it open for future research to theoretically analyze the impact of miscalibration noise on the model.
@@ 4.3 Adversarial Robustness
AD is often used in security-critical areas such as credit card fraud detection or network intrusion detection. Therefore, ML models used for AD must be robust to adversarial attacks. For the following analysis, we note that for low-dimensional, tabular data sets, it is possible that a sample can be completely and effectively transformed into a sample of a different class at high attack strengths. In these cases, the effect on the AUC may be exaggerated and should be interpreted only as an upper bound on the performance drop.
4.3.1 Noise-Free Adversarial Attacks
@@
First, we consider the noise-free model and plot the obtained results of the PGD attacks up to a strength of $\varepsilon=0.5$ in Figure 4. The noise-free QSVR is highly vulnerable to adversarial attacks, as evidenced by the decrease in AUC for small attack strengths of $\varepsilon=0.01$ for all datasets except Toy. The largest decrease occurs for the DoH dataset, whose AUC drops by an order of magnitude from 0.67 to 0.06 for the $\varepsilon=0.01$ attack. As the attack strength increases, the AUC continues to decrease for all datasets until it approaches 0.0 for $p=0.3$. We conclude that techniques to increase the adversarial robustness of the model should be investigated.
@@
The unchanged AUC of Toy for $\varepsilon=0.01$ can be explained by the creation process of the dataset. The dataset was created to be linearly separable with a separation distance of 0.4 between the classes, so the data points must be shifted a large distance in feature space to be misclassified. However, since the AUC drops to 0.0 for $\varepsilon=0.3$, which is smaller than the separation distance, we conclude that the QSVR is susceptible to overfitting.
4.3.2 Noisy Adversarial Attacks
@@
Second, the performance of the noisy QSVR when subjected to adversarial perturbations of strength $\varepsilon=0.1$ is shown in Figure 3. Omitting miscalibration, we find that for noise levels below $p=0.1$, the AUC is low for most noise types and datasets, following the trend of high adversarial vulnerability observed in Figure 4. At higher noise levels, however, the AUC typically recovers to some extent, reaching a plateau at about 0.5. This indicates that the model transitions to a random classifier, showing that the adversarial attacks are so powerful that quantum noise cannot improve performance beyond that of a random classifier. Other researchers report similar results, noting that random noise and adversarial noise are fundamentally different, and that even models resilient to random noise are often still vulnerable to adversarial noise [fawzi2016robustnessclassifiersadversarialrandom].
Miscalibration noise influences both attacked and unattacked models in a similar manner. For models under attack, the plateaus occur at lower AUC values, while near $p=k\pi$ for $k=0,1,2$ the AUC often rises above the adjacent plateaus, producing spikes rather than drops. Such spikes exceed both the AUC at $p=0$ and the 0.5 threshold only for DoH, FMNIST, KDD, Toy, Census, and Mammo., indicating that small amounts of this noise can enhance performance under attack. As the remaining datasets do not exhibit such gains, we conclude that miscalibration noise cannot be regarded as a reliable means of enhancing adversarial robustness.
@@
The DoH dataset is an outlier, with an AUC of 0.0 across almost all noise types and strengths, which is explained by its extreme vulnerability to adversarial attacks seen in Figure 4. This vulnerability can be explained by Figure 5 in the appendix, showing the minimum p-value from the Kolmogorov–Smirnov test and the maximum feature variance within the test set. The DoH dataset has a medium p-value ($\approx 10^{-8}$) combined with a very low variance ($\approx 2\times 10^{-2}$). The p-value indicates the probability that the normal and anomalous samples originate from the same distribution, while the very low variance suggests a high degree of similarity between all samples, especially between the normal and anomalous data. As a result, the DoH dataset is difficult to classify, and even tiny adversarial attacks of strength $\varepsilon=0.01$ lead to manipulations about five times greater than the variance within the dataset.
@@
Notably, the AUC for the DoH dataset with miscalibration noise is $0.0$ across nearly all noise levels and peaks only around $p=\pi$, where it approaches $1.0$. An analysis of the adversarial test kernels for DoH under miscalibration noise (Table 3) highlights major differences between a high-performing run ($p=2.9\approx 0.9\pi$, $\text{AUC}=0.98$) and a low-performing run ($p=1.7\approx 0.5\pi$, $\text{AUC}=0.00$). The mean kernel values in the high-AUC case are four orders of magnitude larger than in the low-AUC case, and hence the disparity between classes is more pronounced, enabling the SVR to separate them more effectively. Since the kernel entries represent the fidelity of the states created by embedding two samples, we conclude that miscalibration noise close to $\pi$ increases within-class similarity and thus elevates the kernel values.
@@
In light of this, the overrotation introduced by miscalibration noise can be interpreted as the addition of fixed-parameter rotation gates independent of the input data. As these parameters strongly affect model performance, this underscores the importance of employing dataset-specific kernels, such as trainable kernels.
@@
Table 3: Analysis of the adversarial test kernel for the DoH dataset subject to miscalibration noise.
@@
4.3.3 Adversarial Training
Adversarial training is a straightforward approach to increasing the adversarial robustness of supervised learning algorithms. Table 4 reveals that adversarial training increases the AUC for the adversarial test set on seven out of eleven datasets, and the average AUC over all datasets rises from 0.28 to 0.31. However, the increase in AUC is small, and except for FMNIST and Toy, the AUC remains below 0.5. For the test set without adversarial samples, the AUC decreases through adversarial training on six out of eleven datasets, and the average declines from 0.75 to 0.71. Table 5 shows the ratio of correctly classified normal samples to the total number of normal samples, as well as the same ratio for the anomalies. For normal data, the ratio is $\frac{\text{tn}}{\text{tn}+\text{fp}}$ , and for the anomalies it is $\frac{\text{tp}}{\text{tp}+\text{fn}}$. We observe that retraining increases the classification ratio for the normal data of the Toy dataset from 0.78 to 0.94, while the ratio for the anomalies remains unchanged at 1.00. For KDD, we report similar results, but the increase in the classification ratio of the normal data through retraining is smaller. This shows that adversarial retraining can lead to more normal samples being classified correctly without influencing the classification of the anomalies, since the latter are not contained in the training set. However, this was not observed for other datasets and was most pronounced for the synthetic dataset, suggesting this behavior requires a large separation distance between the two classes.
@@ 4.3.3 Adversarial Training
Table 4: AUCs of the models for the test set and adversarial test set with and without adversarial training.
@@
Table 5: Ratios of correctly classified adversarial normal and anomalous samples.
@@ 5 CONCLUSION AND OUTLOOK
Second, the influence of six noise channels on the performance of the QSVR was evaluated, revealing that it is largely robust against deporarizing, phasedamping, phase flip and bit flip noise. Amplitude damping, on the other hand, results in the most significant degradation of the model and miscalibration noise also impacts performance.
Finally, the adversarial robustness of the (noisy) QSVR was assessed, and it was observed that the model is highly vulnerable to adversarial attacks. Even weak PGD attacks with a strength of $\varepsilon=0.01$ can reduce the AUC by up to an order of magnitude. Introducing quantum noise does not yield any beneficial effect, neither on the unattacked model’s performance nor on its adversarial robustness. Moreover, adversarial training does not reliably improve the adversarial robustness of the model. Consequently, we conclude that the QSVR demonstrates potential for semisupervised AD in the NISQ era, however, special attention should be paid to the vulnerability to adversarial attacks and amplitude damping and miscalibration noise.
We emphasize the importance of employing dataset-specific kernels and recommend exploring trainable kernels to further enhance the performance of QML models. In addition, a systematic study is needed to understand why, on certain datasets, the hardware-executed QSVR surpasses its simulated counterpart. Furthermore, a theoretical analysis of the effects of miscalibration noise and its influence on kernel values and reconstruction errors will also be helpful. Beyond these directions, future work should explore scaling the model to larger numbers of qubits and, critically, developing effective defenses against adversarial attacks to strengthen the robustness of QSVR in security-sensitive applications.
ACKNOWLEDGEMENTS
@@
APPENDIX
Table 6: Overview of the datasets used for the experiments.
@@
$C$ [$c$] Normal Anomalous
Census [census_income] $\leq$50k >>50k